Poehling Medical Center has a single operating room that is used by local physicians to perform surgical procedures. The cost of using the operating room is accumulated by each patient procedure and includes the direct materials costs (drugs and medical devices), physician surgical time, and operating room overhead. On January 1 of the current year, the annual operating room overhead is estimated to be:

Disposable supplies $187,900
Depreciation expense 33,900
Utilities 19,700
Nurse salaries 282,100
Technician wages 92,400
Total operating room overhead $616,000

The overhead costs will be assigned to procedures based on the number of surgical room hours. Poehling Medical Center expects to use the operating room an average of eight hours per day, seven days per week. In addition, the operating room will be shut down two weeks per year for general repairs.

a. Calculate the estimated number of operating room hours for the year.
hours

b. Determine the predetermined operating room overhead rate for the year.
$ per hour

c. Bill Harris had a 3-hour procedure on January 22. How much operating room overhead would be charged to his procedure, using the rate determined in part (b)?
$

d. During January, the operating room was used 194 hours. The actual overhead costs incurred for January were $41,700. Determine the overapplied operating overhead or underapplied operating overhead for the period. Enter your answer as a positive number.
$  
Answer #1

a. estimated number of operating room hours = Operating hours Per Day * Operating Days in Week * (Operating weeks - Repair Weeks)

Estimated number of operating room hours = 8 * 7 * (52 - 2)

Estimated number of operating room hours = 2800 Hours

b. Predetermined operating room overhead rate = Total Overhead / Operating hours

Predetermined operating room overhead rate = 616000 / 2800

Predetermined operating room overhead rate = $220 per Hour

c. Operating room overhead = Predetermined OH rate * Hours used = $220 * 3 = $660

d. Applied OH = Operating hours * Predetermined OH rate = $220 * 194 = $42680

Overapplied Overhead = Applied Overhead - Actual Overhead = $42680 - $41700 = $980